Marks: --/1
A piece of wire 48 m long is bent into the shape of a rectangle whose length is twice its width. Find the length of the rectangle.
Choose one answer.
a. 8 m
b. 16 m
c. 24 m
d. 32 m


Answer by ewatrrr(24785)


Hi
Let x represent the width. Then 2x would represent the length.
x + x + 2x +2x = 48m
6x = 48m
x = 8m is the width, 16m is the length
checking our answer
2*8m + 2*16m = 16m + 32m = 48m
answer b would be the correct choice